OpenStudy (anonymous):

Write an equation for a line paralell to 6x-5y=15 and travels through the point (5,-2) ?

OpenStudy (***[isuru]***):

hi, When a line in form of "y = mx + c" has an another equation which is parallel to it the gradient of both line should be equal So u have to find the gradient of the equation 6x - 5y = 15 let's say that value is "D" then u can write another equation Y = Dx + C Now u can apply the given co-ordinates in the equation u formed and find C Now write the full equation substituting that value Y =Dx +c and u will have ur answer Can u do that ?

OpenStudy (anonymous):

no im confused

OpenStudy (anonymous):

@jim_thompson5910 mind helping me on this problem please?

OpenStudy (anonymous):

Have you ever plotted a graph? Try to plot y = 2x + 1 and seeing what slope 2 means. For every 2 steps up you go one step right.

OpenStudy (anonymous):

why would i plot y=2x+1?

OpenStudy (anonymous):

Just a suggestion so you can understand the question. Math is not a spectator sport.

OpenStudy (anonymous):

Do you see why y = 2x + 1 and y = 2x + 2 are parallel?

OpenStudy (***[isuru]***):

Alright, I hope u understand wt the gradient is ... gradient is simply the slope of a straight line when there r 2 lines which r parallel to each other this "slope" should be equal and the gradient of a function is represent by "m" in the form Y = mx +c In ur case 6x-5y=15 5y = 6x - 15 y = (6/5)x -3 m= 6/5 so the gradient of the given line is 6/5 Now the equation of the other line also should contain the gradient 6/5 i.e. if Y = mx + C the equation asked in the problem should be somewhat like this y =(6/5)x +C here C represents the intercept of the graph we can find out C by the given co-ordinate (5,-2) i.e. if we substitute that values in the function y =(6/5)x +C we get the function wr the given conditions r satisfied -2 = 6 + c c = -8 so the function of the line that passes through the point C is y = (6/5)x -8 hope this will help u out!

OpenStudy (anonymous):

Wait! I got it, I got y=6/5x+3 :)